The following separate income statements are for Burks Company and its 80 percent–owned subsidiary, Foreman Company:

Burks Foreman
Revenue $ (402,000) $ (302,000)
Expenses 290,000 226,000
Gain on Sale of equipment 0 (16,000)
Equity earnings of subsidiary (56,000) 0
Net Income $ (165,000) $ (92,000)
Outstanding common shares 50,000 40,000

    

Additional Information

  • Amortization expense resulting from Foreman’s excess acquisition-date fair value is $22,000 per year.

  • Burks has convertible preferred stock outstanding. Each of these 6,000 shares is paid a dividend of $5 per year. Each share can be converted into five shares of common stock.

  • Stock warrants to buy 10,000 shares of Foreman are also outstanding. For $15, each warrant can be converted into a share of Foreman’s common stock. The fair value of this stock is $20 throughout the year. Burks owns none of these warrants.

  • Foreman has convertible bonds payable that paid interest of $31,000 (after taxes) during the year. These bonds can be exchanged for 13,000 shares of common stock. Burks holds 10 percent of these bonds, which it bought at book value directly from Foreman.

    Required:

A. Compute basic EPS for Burks Company

B. Compute diluted EPS for Burks Company
Answer #1

Ans: - In this question, we need to calculate basic EPS and diluted EPS for Burks company.

Earning per share is given by(Net income - Preferred Dividends)/ Average of common shares outstanding.

(a) Basic EPS for Burks company.

Particulars Amount
Net Income of Burks company will be (165000-56000) 109000
Adding Net income of Foreman company( 92000-22000) * 0.80 56000
Now Subtract Preferred stock dividend i.e 6000*5 (30000)
Total earnings applicable for Basic EPS for Burks company is 135000
outstanding common shares for Burks company is 50000
Now Basic EPS will be 119000/50000 $2.70

(b) Diluted EPS for Burks company

Particulars Amount.
Net Income for Burks Company will be (165000-56000) 109000
Adding Foreman's share in net income ( see working notes) 67739
Now the earning applicable is 176739
outstanding shares for diluted EPS will be 80000
Diluted EPS for Burks company will be 177439/80000 $2.20

working notes:- Calculating the foreman's earings that is applicable to diluted EPS.

Net income for Foremans company is (92000 - 22000 amortization) = 70000.

Percentage of ownership will be 30000/31000 = 96.77%

Therefore the earnings applicable to diluted EPS will be 96.77% of 70000 = 68439.

Now the outstanding shares for diluted EPS will be given by

Outstanding common shares for Burks company + convertible Preferred stock

50000 + 30000 = 80000.